yhowski Inc. reported income from operations, before taxes, for 2015-2017 as follows:

2015 $173,000
2016 198,000
2017 232,000

When calculating income, Wyhowski deducted depreciation on plant equipment. The equipment was purchased January 1, 2015, at a cost of $72,000. The equipment is expected to last three years and have a(n) $6,000 salvage value. Wyhowski uses straight-line depreciation for book purposes. For tax purposes, depreciation on the equipment is $41,000 in 2015, $17,000 in 2016, and $8,000 in 2017. Wyhowski's tax rate is 35%.

Required:

Enter all amounts as positive numbers.

1. How much did Wyhowski pay in income tax each year? If required, round all calculations to the nearest dollar.

Year Taxes Paid
2015 $
2016 $
2017 $

2. How much income tax expense did Wyhowski record each year?

Year Income Tax Expense
2015 $
2016 $
2017 $

3. What is the balance in the Deferred Income Tax account at the end of 2015, 2016, and 2017? If your answer is zero, enter "0". If required, round all calculations to the nearest dollar.

Year Balance Debit or Credit
2015 $ Credit
2016 $ Credit
2017 $ No balance
Answer #1

Solution (1)

Lets calculate income from operations for tax purposes-

Recorded Profits Given

2015 $173,000
2016 198,000
2017 232,000

As per the question company follows straight line depriciation for book purposes. Equipment is purchased for $72,000 life 3 years having a salvage value of $6,000

so Yearly depreciation = actual cost - salvage / life

= 72,000 - 6,000= $22,000 / year

Difference between Book Depreciation and depreciation for tax purposes-

YEAR 2015 = 41,000-22,000 = 19,000 $

YEAR 2016 = 17,000-22,000 = - 5,000 $

YEAR 2017 = 8,000-22,000 = - 14,000 $

lets calculate income tax paid

tax paid in year 2015 = (income from operations- difference between book depriciation and depriciation for tax purposes) x tax rate

= (173,000- 19,000) x35% = $ 53,900

tax paid in year 2016 = (income from operations- difference between book depriciation and depriciation for tax purposes) x tax rate

= [198,000-(-5,000)] x 35% = $ 71,050

tax paid in year 2017= (income from operations- difference between book depriciation and depriciation for tax purposes) x tax rate

= [232,000-(-14,000)] x 35% = $86,100

Solution (2)

lets calculate income tax expense Wyhowski record each year in books-

Tax expenses in year 2015 = 173,000 x 35 % = $ 60,550

tax expenses in year 2016 = 198,000 x 35% = $ 69,300

tax expenses in year 2017 = 232,000 x 35 % = $ 81,200

Solution (3)

the balances in the Deferred Income Tax account

2015 = $ 6,650 Credit

2016 = $ 4,900 Credit

2017 = $ 0